[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon May 22, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Mon May 22 12:40:09 UTC 2017


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
There were problems sending out several products from AWIPS Sunday evening  
and Monday morning. This was due to some bad disc drives on NHC-N. NCF  
reconfigured the system to use backup drives, and Chris Mello is  
coordinating to get new drives shipped to us ASAP.
HSU:
Himawari data is missing from N-AWIPS. Monica is investigating.

The Meteosat data is still missing over the Indian Ocean region. That  
happened when Met7 was replaced by Met8 a few months ago. TSB will need to  
modify our satellite ingest systems to get Met8 instead of Met7. It will  
probably be a few more weeks before we can get to this since we have higher  
priority issues ahead of this.

The seahorse3 server that shares with worksheet with the WFOs and hosts  
some model field displays had a disk problem on Friday. NCO has been  
notified and we expect them to address this today. If there was an active  
storm, we would have called in off-hour support.
Other items:
We are still waiting for NCO to implement our higher bandwidth so the GFS  
and other models come in sooner. NCO said this is about 2 weeks away.

An AWIPS patch will be installed at about 10:30 AM today, which will  
require about 1 hour of down time. TAFB has been notified and will work  
around the outage. Chris Mello is coordinating this upgrade with NCF.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170522/90428351/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list